OpenStudy (anonymous):

Using the quadratic formula to solve 11^2 – 4x = 1, what are the values of x?

OpenStudy (anonymous):

Make the equation equal to 0 first.

OpenStudy (anonymous):

@dreamstoogood could you double check the equation? thanks

OpenStudy (anonymous):

A. \[\frac{ 2 }{ 11 }\pm \frac{ \sqrt{15} }{ 11}\] B. \[\frac{ 2 }{ 11 } \pm \frac{ 2\sqrt{15} }{ 11}\] C.\[\frac{ 2 }{ 11}\pm \frac{ \sqrt{7} }{ 11 }\] D.\[\frac{ 2 }{ 11 }\pm \frac{ \sqrt{7}i }{ 11 }\]

OpenStudy (anonymous):

I assume that's 11x^2-4x=1